OpenStudy (anonymous):

what is the weight in lbs of 1 cu ft cork if weighs 0.25 gm.

OpenStudy (compphysgeek):

the conversion factor is \[\frac{1 lb}{453.6 grams}\] that should be all you need

OpenStudy (anonymous):

cu means cubic?

OpenStudy (compphysgeek):

yes, that's right. and a cubic feet of cork weighs 0.25 grams or did I get it wrong?

OpenStudy (anonymous):

i need a formula .. bcuz i dont have any idea how to do that

OpenStudy (compphysgeek):

if it's really 0.25 grams than the formula is similar to all the others that we used so far \[ 0.25\,g \times \frac{1 lb}{453.6\,g} = 5.51×10^{-4}\,lb \] that seems to be very light

OpenStudy (anonymous):

thats the formula?

OpenStudy (compphysgeek):

unless I got the information wrong. In the formular to start with what you've got, grams in this case, you multiply with the conversion factor, the grams cancel out and you're left with pounds, what you were looking for.

OpenStudy (anonymous):

what is the weight in lbs of 1 cu of a cork

OpenStudy (compphysgeek):

is there more information? is the cork's mass 0.25 grams or is it 0.25 grams/ cubic centimeter?

OpenStudy (anonymous):

this the q. what is the weight in lbs of 1 cu ft of cork if on weighs 0.25gm.?

OpenStudy (compphysgeek):

and gm is grams? if so then all you need is what I posted above.

OpenStudy (anonymous):

ok i get the answer

OpenStudy (compphysgeek):

ok